EVA Report – February 27th

Crew 206 EVA Report 27-Feb-2019
EVA #3

Author: Aurélien Mure

Purpose of EVA: The EVA is built to begin the exploration and to
familiarize the crew with mid range EVA’s. We will go with the rovers to
the moons region and start hiking from this point. The purpose will be
to explore by walking The Moon, Gray Moon, Beige Moon and Yellow Moon.
The rovers will be parked at the intersection between Copernicus Hwy and
Brahe Hwy, the rest will be done by walking counterclockwise starting
with Yellow Moon. It is important for us to start doing long EVA’s to be
used to it quickly. As 2 of us will be commanders of crews in 2020 and
we can’t know yet who would it be, the whole group needs to be trained
to all types of EVA’s. That is why we will continue to ask for EVA’s
with four marsonauts. Moreover, in our mind being two on each rover is
safer than being alone.

If we realize that the batteries won’t allow this, we will stop at White
moon and explore by hiking Beige Moon and Gray Moon, or if we have to
stop before near Reservoir Dam and Cowboy Corner.

Start time: 9:08 am

End time: 11:14 am

Narrative: After the 5 minute depressurization, our team Blue (Gaspard
and me) checked the rovers state, they were charged and ready. Team Red
(Benjamin and Norbert) checked the water level of the static tank. Team
Blue will ride Opportunity and Team Red Curiosity. The departure was at
9:18. We drove to the LOAC place to check quickly the battery level.
12,2V, everything is nominal, our experiment is on its ay. We drove
carefully with as objective saving energy. We checked the battery every
2 min to decide if we went to our plan A (Copernicus Hwy), plan B
(beginning of Brahe Hwy) or plan C (Reservoir Dam). We went through our
plan C destination with enough energy to continue until plan B
destination. Just after Reservoir Dam we decided to focus on plan B. We
arrived at the beginning of Brahe Hwy with 67% of battery for
Opportunity and 72% for Curiosity. We went for a walk in the region of
Grey Moon, for an hour and around 3km. The ground was very slippery so we
were extremely careful with our steps.

At 10:46 we decided to leave the place and go back to the Hab. We
arrived at 11:06, with Curiosity at 61% and Opportunity at 34%. With
Spirit that is supposed to be the laziest one, we stick to the reality
as on mars Curiosity is in operation, Opportunity has died 4 months ago
and Spirit a long time ago !

Destination: Gray Moon
Coordinates : 516450/4254800

Participants: Aurélien Mure (XO, EVA leader), Benjamin Auzou (COMMS ),
Gaspard Thieulin (ENG), Norbert Pouzin (GHO)
Road(s) and routes per MDRS Map : Entrance Road , Cow Dung Road

Mode of travel: Driving and Walking

Greenhab Report – February 27th

Crew 206 Green Hab Report – 27-FEB-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 18.2°C

High: 35.9°C

Hours of supplemental light: 7 Hours

Daily water usage for crops: 8.6 gallons

Water in Blue Tank: 276.5 gallons

Crops watered at 0940 hours (2.2 gallons) and 1830 hours (6.4 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 09.40 AM 35.0° C / 0 %
2) 11.30 AM 24.8° C / 0 %
3) 15:00 PM 21.2° C / 24 %
4) 18:30 PM 22.8° C / 33 %

Changes to crops: I replanted radishes.
Radishes from my experiment are starting sproutling.
Carrots replanted yesterday don’t all look great.
Tomatoes look great, but in two pots they haven’t germinated.

Narrative:
This morning when Jeremy took care of the greenhab (I was in EVA), it
was really hot. We had difficulties to monitor the heat of the hab in
the morning : the fan cooled too much the greenhab.
The door was opened from 0940 to 1630 ; the fan from 1100 to 1130.
I replanted some radishes.

Harvest: None.

Support/supplies needed: None.

Astronomy Report – February 27th

Astronomy Report, Sol 3
Name: Aurélien Mure
Crew: 206
Date: 02/25/2019
ROBOTIC OBSERVATORY
Object observed : Orion Nebula, M42. I attach the image I did from the observation. I’m a beginner in astrophotography, I hope my pictures will be improved during the mission.
Problems encountered : I think I exposed to much the telecope, the center of the nebula is to bright.

Commander Report – November 25th

Astronomy Report, Sol 2

Name: Aurélien Mure

Crew: 206
Date: 02/25/2019

MUSK OBSERVATORY

Solar Features Observed: None, no clouds visible but I was unable to focus on features, maybe because of the very low sun activtiy, tell me if I’m wrong.

Problems encountered : How do I move the telescope to be faced to the sun when the alignment is set ?

Hi Eric and Peter,
It was a long day of astronomy for me, I observed the sun, tried to work on my images of the crab nebula and planned the observation of M42 for tonight.
I have some questions to ask you :
– How can I use the hand control to move the solar telescope when it is focused on the sun and not well aligned ?
– Is it still normal that I can’t see any solar feature with any configuration of the telescope ?
I hope my first images will be great !
Thanks,
Aurélien

Journalist Report – February 26th

Crew 206 – 02/26/2019

Sol 2

Author : Auzou Benjamin, Journalist

“La Vie En Rouge”

At the precise moment my feet touched the Martian ground, and my eyes raised up to the sky, the time was liked stopped in front of me. There are incredible places that challenge our imagination, with their shapes, their reliefs, their colors. The Martian desert I had the chance to admire this morning is one of those. In front of this marvelous theater, I wish I had the talent of Picasso, Monet or Turner to capture
that precise moment, that sky, that landscape only a painting can give back the magnificence.

But this EVA had above all a scientific goal: deploy the LOAC (an aerosol counter) and the weather station. It was my first EVA, along with Cerise, Jérémy and Gaspard, on the rovers Curiosity and Opportunity. Quite funny that not so long ago, in 2018, they were the only two active “beings” at the surface of Mars.

Aurélien, HabCom of the day, helped us to spend the five minutes of the depressurization with a musical ambiance: La Vie en Rose by Edith Piaf. The words to describe this day came directly to me: La Vie en Rouge.
After the concentration and the rigor of the EVA, a surprise were waiting for us inside: a pizza crust prepared by Aurélien, while Norbert was bringing his harvest : basil, sage and parsley, in order to give a providential freshness to the meal. We shared two pizzas that we named Verde and Roja. Aurélien and Norbert logically take the head of the Martian food contest.

Then the afternoon was the first one of a routine that will last all through the mission with both human factors experiments: rover driving and virtual reality.

The Sun is falling down on Mars, giving its last colors to the landscape, before putting it in obscurity and let the stars shine in the pure sky. I glance at the horizon, keeping that moment in my memory for ever, today, I walked on Mars.

Greenhab Report – February 26th

Crew 206 Green Hab Report – 26-FEB-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 15.9°C

High: 28.5°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 8.6 gallons

Water in Blue Tank: ~ 285.1 gallons

Crops watered at 0905 hours (2.2 gallons) and 1830 hours (6.4 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 09.05 AM 24.0° C / 26 %
2) 11.30 AM 24.2° C / 0 %
3) 15:00 PM 23.1° C / 0 %
4) 18:30 PM 21.6° C / 46 %

Changes to crops: I replanted some carrots, and a read leaf lettuce. I pulled out flowering greens.
Spinach and mustard, in the pots, are sprouting. Peas in the shelf still look bad (quite yellow) even with the daily use of Miracle Gro.

Narrative:
This morning I replanted some carrots from the blue gardener to pots, so they have more place to grow. A lot of carrots needs to be replanted, the blue gardener is quite full. I will do it in the next days. I also replanted a lettuce.
This afternoon I pulled up the flowering carrots and radishes. Total biomass weight : 22 g of carrots and 17 g of radishes.

To monitor the heat in the hab, I opened the door at 0900, and closed it at 1630.
At 1100 the heat increased (28°C), so I started the fan (from 1100 to 1130 and then from 1430 to 1500) to cool down the greenhab. It worked well, since the heat in the afternoon didn’t reached more than 26°C.

As asked, I changed the hours of supplemental light : the light is now on from 0500 to 0730, and from 1730 to 2200. So we will have 7 hours of supplemental light tomorrow. Tell me if it is too much !

I harvested lots of basil, parsley, and sage for a homemade pizza, by our chief Aurélien !

Harvest: 11.2 g basil, 3.0 g of lemon basil, 10.4 g of parsley and 1.4 g of sage. 39 g of Biomass (flowering plants).

Support/supplies needed: None.

Journalist Report – February 26th

Crew 206 – 02/26/2019

Sol 2

Author : Auzou Benjamin, Journalist

“La Vie En Rouge”

At the precise moment my feet touched the Martian ground, and my eyes raised up to the sky, the time was liked stopped in front of me. There are incredible places that challenge our imagination, with their shapes, their reliefs, their colors. The Martian desert I had the chance to admire this morning is one of those. In front of this marvelous theater, I the talent of Picasso, Monet or Turner to capture that precise moment, that sky, that landscape only a painting can give back the magnificence.

But this EVA had above all a scientific goal: deploy the LOAC (an aerosol counter) and the weather station. It was my first EVA, along with Cerise, Jérémy and Gaspard, on the rovers Curiosity and Opportunity. Quite funny that not so long ago, in 2018, they were the only two active “beings” at the surface of Mars.

Aurélien, HabCom of the day, helped us to spend the five minutes of the depressurization with a musical ambiance: La Vie en Rose by Edith Piaf. The words to describe this day came directly to me: La Vie en Rouge.
After the concentration and the rigor of the EVA, a surprise were waiting for us inside: a pizza crust prepared by Aurélien, while Norbert was bringing his harvest : basil, soy and parsley, in order to give a providential freshness to the meal. We shared two pizzas that we named Verde and Roja. Aurélien and Norbert logically take the head of the Martian food contest.

Then the afternoon was the first one of a routine that will last all through the mission with both human factors experiments: rover driving and virtual reality.

The Sun is falling down on Mars, giving its last colors to the landscape, before putting it in obscurity and let the stars shine in the pure sky. I glance at the horizon, keeping that moment in my memory for ever, today, I walked on Mars.

Sol Summary – February 26th

Crew 206 February 26th 2019
Sol 2

Summary Title : Busy Day

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Today we slept half an hour more than yesterday, we woke at 07:00 am. After our 30 minutes workout, we ate quickly to have enough time to prepare our EVA. We had to gather all the necessary tools and organise the different pieces of the weather station and the LOAC. Around 09:00 am, we were almost ready. We left the airlock at 09:17 am, and took the rovers to carry our equipment to the site we had found on EVA #1. The installation went fine, even though we had to go back to the Hab twice to get additional tools. The two instruments are installed, but we still have to make some adjustments. We finished the EVA at 10:30 am. After that we worked on our experiments and our reports. For lunch Aurélien and Norbert made some pizzas, they were great. The first main activity of the afternoon was TELEOP, the human factor experiment. We had an issue with the tele-operated rover, so today we only worked with the simulated rover. This part of the experiment worked fine. After that we participated in the Virtual Reality experiment once again.

Look ahead plan : Tomorrow we will start working on the simulation space suit, and it will also be a little rest day for Cerise and I that did two EVAs in a row. We will follow up our work on our various experiments.

Anomalies in work : none

Weather : Quite sunny all day, just a few small clouds

Crew physical status : small aches from physical training but fine otherwise

EVA : 500 meters from the Hab (North East), 40 meters from Cow Dung road. Installation of the LOAC and the weather station

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA Request

Support Requested : a new frying pan, a new sponge for the dishes, a rag and a tissue napkin to prevent using too much paper towels

Operations Report – February 26th

Crew 206 Operations Report 26 – FEB – 2019

SOL: 2

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Generator (hours run): Turned off at 0800 (24th Feb), Turned on at 1800 (24th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 82%, Turned on Gen (24th Feb) 64%

Diesel Reading – ~ 75%

Propane Reading – ~ 23%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 93%, ~ 509 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 285,1 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01415823 units

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Opportunity rover used: Used
Hours: 51.7 hours
Beginning charge: 100%
Ending charge: 95%
Currently charging: Yes

Curiosity rover used: Used
Hours: 75.9 hours
Beginning charge: 100%
Ending charge: 93%
Currently charging: Yes

Spirit rover used: Not Used
Hours: 78.3 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Notes on rovers: Charging socket has been fixed, everything is currently fine.

ATV’s Used: N/A

Reason for use: Not used
Oil Added? N/A
ATV Fuel Used: N/A
# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used

Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Lost Internet access once in the day. Came back after no more than 10 minutes. Probably the same issue as yesterday.

Summary of suits and radios: Checked all the suits batteries. They are all fine and charging. Except for suit #11 on which the charging socket is disconnected from the battery. Will fix it ASAP. Checked suit #1 (had fogging issues yesterday) fans : they are working fine. Maybe the battery wasn’t charged enough for EVA #1.

Summary of Hab operations: Added some tape to the inside locker for engineering airlock. The door now closes more precisely.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations:

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.